Draw all possible samples of size 2 with replacement from population :  2  4   6  8   10.

prove that population mean is same as samples mean.


1
Expert's answer
2021-09-27T07:39:00-0400

In total there are "5^2=25" samples or pairs which are possible

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 Sample & Mean & & Sample & Mean & & \\ \hline
 (2,2) & 2 & & (2,4) & 3 \\
 \hdashline
 (2,6) & 4 & & (2,8) & 5 \\
 \hdashline
 (2,10) & 6 & & (4,2) & 3 \\
 \hdashline
 (4,4) & 4 & & (4,6) & 5 \\
 \hdashline
 (4,8) & 6 & & (4,10) & 7 \\
 \hdashline
 (6,2) & 4 & & (6,4) & 5 \\
 \hdashline
 (6,6) & 6 & & (6,8) & 7 \\
 \hdashline
 (6,10) & 8 & & (8,2) & 5 \\
 \hdashline
 (8,4) & 6 & & (8,6) & 7 \\
 \hdashline
 (8,8) & 8 & & (8,10) & 9 \\
 \hdashline
 (10,2) & 6 & & (10,4) & 7 \\
 \hdashline
 (10,6) & 8 & & (10,8) & 9 \\
 \hdashline
 (10,10) & 10 & & \\
 
\end{array}"

The table shows that there are nine possible values of the sample mean "\bar{X}."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \bar{x} & p(\bar{x}) \\ \hline
 2 & 1/25 \\
 3 & 2/25 \\
 4 & 3/25 \\
 5 & 4/25 \\
 6 & 5/25 \\
 7 & 4/25 \\
 8 & 3/25 \\
 9 & 2/25 \\
 10 & 1/25 \\
\end{array}"

"\mu_{\bar{X}}=\sum_i\bar{x}_ip(\bar{x}_i)=2(\dfrac{1}{25})+3(\dfrac{2}{25})+4(\dfrac{3}{25})"

"+5(\dfrac{4}{25})+6(\dfrac{5}{25})+7(\dfrac{4}{25})+8(\dfrac{3}{25})+9(\dfrac{2}{25})"

"+10(\dfrac{1}{25})=6"

"\sum_i\bar{x}_i^2p(\bar{x}_i)=(2)^2(\dfrac{1}{25})+(3)^2(\dfrac{2}{25})+(4)^2(\dfrac{3}{25})"

"+(5)^2(\dfrac{4}{25})+(6)^2(\dfrac{5}{25})+(7)^2(\dfrac{4}{25})+(8)^2(\dfrac{3}{25})"

"+(9)^2(\dfrac{2}{25})+(10)^2(\dfrac{1}{25})=40"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\sum_i\bar{x}_i^2p(\bar{x}_i)-(\sum_i\bar{x}_ip(\bar{x}_i))^2"

"=40-(6)^2=4"

"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{100}=10"

"\mu=\dfrac{2+4+6+8+10}{5}=6"

In sampling with replacement the mean of all sample means equals the mean of the population:

"\mu_{\bar{X}}=6=\mu"

"\sigma^2=\dfrac{1}{5}((2-6)^2+(4-6)^2+(6-6)^2"

"+(8-6)^2+(10-6)^2)=8"

When sampling with replacement the variation of all sample means equals the variation of the population divided by the sample size 


"\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}"

"4=\dfrac{8}{2}"